by TEI TEI

Building Operations Project

08/06/15 – Temperature probes put in space to monitor temperatures. Temps and space seemed to be tempered well. Temp and humidity within reason. ThermalTech collected original design criteria from Trane on AH-21 and all the other air handlers ordered at the same time from Trane and shared this with the County.

08/12/15 – Preliminary report sent to Tony Matre. Unit is undersized for space at full load and number of people for space. Trace load calc shows 12.5 to 15 tons of cooling required. Existing unit is 5-7.5 tons but space seemed conditioned much better than this.

08/19/15 – Tony requested new SOSS to replace AH-21.

09/23/15 – Tony Matre asked for an SOSS to replace AH-21.

10/14/15 – We decided to bid this project as a Small Project bid.

10/28/15 – ThermalTech met with Facilities to start filed work today on this project for small project bid.

11/25/15 – All bids came in above $50,000. Bidders were advised that no contract could be awarded. ThermalTech contact Trane immediately to inquire about a rebuild project under State Term we had discussed previously. Trane started working on a new proposal immediately.

12/04/15 – Trane submitted a new price. TEI approved scope and cost for a preliminary PO for this work. New coil and motor should help meet max conditions in this room.

01/12/16 – TEI sent state term contract in for approval with Tony and Don Webber.

02/17/16 – Don Webber sent back and email saying he was awaiting comments from County to proceed with work.

02/26/16 – Tony says contract just needs to be signed by Trane and we can kickoff the job. TTC will advise Trane.

03/30/16 – Contract and PO approved, equipment is ordered, waiting for Trane to start.

04/15/16 – Equipment is ordered and will be here in a few weeks. Trane getting background checks in order.

05/18/16 – Trane technicians on site setting up project work. There is discussion about the motor upgrade and VFD size.

06/22/16 – TTC sent out the punchlist and asked Trane for feedback on why the bigger motor was not installed. TEI also recommend that Andy increase the SP of AH-21 to 1.5″.

06/30/16 – Trane received the punchlist and thought they many not have add the motor to make the unit work with the the new coil. They are verifying this. TTC dropped off 5 temp sensors on Tuesday to track temperatures in space. Apparently the worst time is coming out of nigh setback on really hot days. TTC and Tony discussed the next steps after the VFD is up and running at 100%.

07/18/16 – TTC sent VFD data to Trane and asked them to continue with the motor replacement as included in the contract. TTC sent summary field report to Andy and Tony for Room 260 showing that the room is not maintaining temperatures when its over 90F outside and cannot reach 72 on Monday until 4:00pm in the afternoon after hot weekends.

07/21/16 – Don Webber from Trane and his technician will be the Courthouse on Friday to look at installing the motor as part of the contract work for AH-21. My fieldwork shows we are not meeting temperatures on over 90F days and coming out of night setback on the weekends.

07/27/16 – Talked to Tony and Andy, Trane was down and checked the motor and VFD and told Tony that something did not match up. I am waiting for more details from Trane as promised.

08/03/16 – TTC check in with Trane on 8/1, no updates yet.

08/11/16 – Trane sent field reports, TTC looked over the reports and responded requesting a phone conference.

08/31/16 – TTC talked to Andy and request more info from Trane. Running the system at 100% still does not cut it on really hot days and hot weekends coming out of nigh setback

09/08/16 – Closeout meeting scheduled with Trane next Friday 9/16.

09/16/16 – ThermalTech met with Trane and Courthouse to discuss next steps. Trane will resheave the fan, ThermalTech will take the air readings next week.

09/27/16 – ThermalTech was on site tonight getting airflow measurements on this unit. TEI to issue field progress report to Tony and Andy. The basic readings were that AH-21 is delivering at least 2500 cfm so it meets the original design.

11/10/16 – We have contacted Trane to resheave the motor again. The extra duct was traced out by Andy and Terry and fees only the hallways, it can probably be throttled back.

11/16/16 – TTC contacted Don Webber again for a status update on this project. He is out until 11/21.

11/25/16 – Trane said that they were sending at technician down to resheave on 11/17.

12/09/16 – Work is complete by Trane and unit is resheaved. TTC and Andy to verify next spring for cooling load capability.